### v2

@@ 1 Introduction
Rabiner introduces HMMs [16] and proposes to use the Forward-Backward algorithm [2] for inference. Zhang et al. distill a language model into an HMM [20]. HMMs can be extended to Dynamic Bayesian Networks [7], which can model interactions between different state variables. Inference is done by, e.g, the interface algorithm [14]. However, query answering in Bayesian networks [15] is NP-hard [6], giving rise to approximations. Boyen and Koller represent the distribution over the random variables with successors in the next time step as a product of marginals [5], which is extended by Murphy and Weiss to always use a factored representation [13]. Instead of instantiating a DBN for each time step, Gao et al. introduce a sliding window approach [8]. Murphy provides an overview over approximate inference [14]. Though, the focus in literature is on approximate inference in DBNs or approximation for learning HMMs, but not on approximate inference within HMMs, which we propose in this paper. Vithanage et al. [19] effectively use the $k$ most probable transitions and show that this minimizes the Kullback-Leibler divergence if restricting to $k$ transitions. Contrary, we provide a more fine-grained error assertion.
@@ 2 Hidden Markov Models Generalizing Dynamic Models
A Hidden Markov Model consists of two series of random variables $(S_{t},O_{t})_{t}$ over a set $S$ of states and $O$ of observations, where $S_{t}$ denotes the state in time step $t$ and $O_{t}$ the respective observation. The probability $P(O_{t}\mid S_{t})$ of an observation depends only on the current state. The probability distribution over the states is given as the temporal behavior $P(S_{t}\mid S_{t-1})$ and a prior $P(S_{0})$. The semantics for a time step $t$ is thus given by
@@
where $S_{t}$ is the distribution specifying the probabilities $s_{t}$ for each state $s\in S$.
@@
Now, the important part is that the sum in Equation 1 goes over all possible states, rendering the equation inefficient for large state spaces, like in language models. Our goal is to reduce the number of states that have to be enumerated. While an HMM seems simple, it can be thought of as a generalized dynamic PGM: We have a, potentially complex, state $S_{i}$ emitting observations $O_{i}$. Let us denote the distributions $P(s_{t}),\,s\in S,$ as forward message updating the state probabilities through time. Dynamic Bayesian Networks (DBNs) [7] can be viewed as HMMs by treating the joint probability distribution over the interface [14] as state and all other random variables jointly as observation. Furthermore, GPT-based language models learn the forward message in a highly complex state space. At their core, they can be viewed as HMMs, interpreting, analogous to DBNs, each possible complex internal state of the language model as a state of an HMM with respective transition and observation (or emission) probabilities. Generating sentences from language models is then the same as predicting states and sampling observations from HMMs, which leads to a noisy distribution in the far future without any observations.
@@ 3 Inference in Top-$p$ Hidden Markov Models
In other words, the probabilities for all events not in the top-$p$ probabilities are set to zero while the other probabilities are scaled to sum up to one. For Example 2 and $p=0.9$, Table 2 shows the top-$p$ distribution. As an HMM consists of conditional probability distributions, we independently apply the top-$p$ distributions to them:
@@
Given an HMM $H$ with a set $S$ of states, a set $O$ of observations, probability distributions over the states given by $P_{H}(S_{0})$ and $P_{H}(S_{i+1}\mid s_{i})$, for each state $s\in S$, and probability distributions $P_{H}(O_{i}\mid s_{i})$, $s\in S$, over the observations, the top-$p$ HMM of $H$ is another HMM $Q$ over the same set $S$ of states and same set $O$ of observations. The probability distributions of $Q$ are the top-$p$ distributions of the respective ones of $H$, that is, (i) $P_{Q}(S_{0})$is the top-$p$ distribution of $P_{H}(S_{0})$, (ii) $P_{Q}(S_{i+1}\mid s_{i})$is the top-$p$ distribution of $P_{Q}(S_{i+1}\mid s_{i})$, for each state $s\in S$, and (iii) $P_{Q}(O_{i}\mid s_{i})$is the top-$p$ distribution of $P_{H}(O_{i}\mid s_{i})$, $s_{i}\in S$.
@@
Having shifted the probability mass, we now exploit the new distributions for efficient inference without enumerating all states. Since the transition matrix still has the same dimensions, but is more spare, we use sparse matrices to not enumerate all states. We explain how to perform inference, by (sparse) matrix multiplication, in a top-$p$ HMM $Q$ of an original HMM: (1) The top-$p$ HMM $Q$ is built according to Definition 4 by using Algorithm 1. (2) The obtained transition probabilities and observation probabilities are stored in a transition and observation matrix, respectively. (3) The two matrices are each transformed into compressed row storage [1]: For each row in the matrix, a tuple $(c,x)$ for each non-zero entry, where $c$ is the column number and $x$ the value. (4) For inference, sparse matrix-vector multiplication [4] is used to proceed the current state distribution in time and to get the observation distribution: The dot product between each row and the vector is calculated by summing the products of $x$ and the entry $c$ of the vector for each $(c,x)$ tuple in the row. Sparse matrix-vector multiplication runs in time proportional to the number of non-zero entries [4]. Sparse matrix-vector multiplication is also available for GPUs [3].
@@ 4.2 Overall Approximation Error
Fixing two state distributions $\varphi$ and $\psi$, the transition of a stochastic process $Q$ can be split into two steps: First, a transition defined by $R^{\Gamma}$ from $\Omega$ to $\tilde{\Omega}$ and, second, a transition defined by $R^{\Delta}$ from $\tilde{\Omega}$ to $\Omega$, where $\Omega=\left\{\omega_{i}\right\}_{i}$ is the state space of $Q$ and $\tilde{\Omega}=\Omega\cup\{c\}$ with a new intermediate state $c$. The process $R^{\Gamma}$ preserves its state with probability $1-\gamma$ and transitions to state $c$ with probability $\gamma$. The process $R^{\Delta}$ transitions from $\omega_{i}$ to $\omega_{j}$ with probability $\frac{Q_{i,j}^{\Delta}}{1-\gamma}$. For state $c$, $R^{\Delta}$ it transitions to state $\omega_{j}$ with probability $\sum_{i}\varphi(\omega_{i})\frac{Q_{i,j}^{\Gamma}}{\gamma}$. The matrices $Q^{\Gamma}$ and $Q^{\Delta}$ arise from an additive contraction decomposition $Q=Q^{\Gamma}+Q^{\Delta}$, where $Q^{\Gamma}$ and $Q^{\Delta}$ are non-negative matrices such that, for all $i$, $\sum_{j}Q_{i,j}^{\Gamma}=\gamma$, and for all $j$, $\sum_{i}\varphi(\omega_{i})Q_{i,j}^{\Gamma}=\sum_{i}\psi(\omega_{i})Q_{i,j}^{\Gamma}$.
@@
$\delta(\varphi^{\prime},\psi^{\prime})\leq\left(1-\gamma_{Q}\right)\delta(\varphi,\psi).$ (14)
@@ 5 Evaluation
We evaluate the top-$p$ approach on three HMMs: One favoring our approach, one that is disadvantageous for the top-$p$ approach and a language model for a realistic application. We investigate the sparsity, the runtime with and without observations, and the total variation distance of the top-$p$ model. We have trained a very simple language model as an HMM with 7620 states, which we call LM HMM. In practice, one would distill a larger model into an HMM, like in Zhang et al. [20]. In this evaluation, we focus on the runtime and not on the quality of the language model. We create a synthetic Bell HMM, which is designed to favor the top-$p$ approach. The Bell HMM distributes, in each next-state distribution, a probability mass of 0.9 over five states and the remaining 0.1 over 795 states, favoring our approach by having lots of unlikely successor states. For a synthetic HMM disadvantageous for the top-$p$ approach, we create a Uniform HMM consisting of uniform state distributions for a system of 800 states. The number of observations equals the number of states in all HMMs. We evaluate top-$p$ for the $p$ values $0.5$, $0.7$, and $0.9$ and for the maximum time step $50$. We mostly give only the numbers for $p=0.9$ as the three top-$p$ approaches are mostly comparable. We run all tests on a 13th Gen Intel(R) Core(TM) i5-1345U CPU with 1.60 GHz and 16 GB of RAM. We use NumPy [10] and SciPy [18] for matrix and sparse matrix operations, respectively.
@@
Sparsity We define the sparsity as the ratio of zero-entries in the transition matrix. In the Bell HMM, all three top-$p$ approaches lead to a sparsity of above $0.99$. In the Uniform HMM, the sparsity is approximately 1p1-p. And in the LM HMM, the sparsity is for all three top-$p$ greater than $0.97$.
@@
Runtime for Inference The top-$0.9$ HMM for the Bell HMM outperforms exact inference in the original HMM by a factor of more than $15$. For the Uniform HMM, top-$0.9$ is around $3.39$ times slower than the original HMM. The reason for the slower runtime is the low sparsity: Sparse matrix multiplication only pays off for high sparsity. In the LM HMM, top-$p$ is more than $18$ times faster than the original HMM. Figure 1(a) shows the experiment for the LM HMM. In Appendix 0.A.1, we show the experiment for the Bell and Uniform HMM.
Runtime for Filtering Before, we reported runtimes without observations. Now, we enter an observation every five time steps. We omit the LM HMM, since observation in text generation is not applicable. In the Bell HMM, the top-$0.9$ approach is more than $77$ times faster than filtering in the original HMM. In the Uniform HMM, the top-$0.9$ approach is more than $24$ times faster than the original HMM. Figure 1(b) shows the experiment for the Bell HMM. In Appendix 0.A.1, we show the experiment for the Uniform HMM.
Total Variation Distance The total variation in the Bell HMM is $0.086$ for top-$0.9$, $0.285$ for top-$0.7$ and $0.78$ for top-$0.5$. In the Uniform HMM, the total variation distance is $0.099$, $0.299$ and $0.499$ for top-$0.9$, top-$0.7$ and top-$0.5$, respectively. In the LM HMM, the total variation distance is $0.059$, $0.172$ and $0.3$ for top-$0.9$, top-$0.7$ and top-$0.5$, respectively.
Total Variation Distance for Filtering With observations, the total variation in the Bell HMM is $0.095$ for top-$0.9$, $0.491$ for top-$0.7$ and $0.846$ for top-$0.5$. In the Uniform HMM, the total variation distance is the same as for no observations, since the model is uniform. We again omit the LM HMM, since observation in text generation is not applicable.
(a) LM HMM
(b) Bell HMM with filtering
@@
Figure 1: Runtime for inference in the LM and Bell HMMs. The blue line shows the runtime in the original HMM and the other three ones for different top-$p$ ones.
@@
Summing up, the top-$p$ approach significantly increases sparsity to more than $0.9$ for top-$0.9$. Moreover, top-$p$ is at least an order of magnitude faster than the original HMMs, while the total variation distance for top-$0.9$ stays below $0.09$.
@@ 6.0.1 Acknowledgements
Appendix 0.A Appendix
0.A.1 Evaluation
In this appendix, we show more plots for the evaluation in Section 5.
Runtime for Inference
Figure 2 shows the runtime for inference in the Bell HMMs. All three top-$p$ approaches are significantly faster than inference in the original HMM and are comparable fast. Figure 3 shows the runtime for inference in the Uniform HMM. Clearly, the lower $p$, the faster the inference. However, all top-$p$ models have a slower runtime than the original one. Both is because the sparsity increases with higher $p$ and sparse matrix multiplication only pays off when the sparsity is high.
Figure 2: Runtime for inference in the Bell HMM. The blue line shows the runtime in the original HMM and the other three ones for different top-$p$ ones.
Figure 3: Runtime for inference in the Uniform HMM. The blue line shows the runtime in the original HMM and the other three ones for different top-$p$ ones.
Runtime for Filtering
Figure 4 shows the runtime for filtering every five steps. In contrast to regular prediction, the top-$p$ approaches are now significantly faster than filtering in the original HMM.
Figure 4: Runtime for filtering in the Uniform HMM. The blue line shows the runtime in the original HMM and the other three ones for different top-$p$ ones.

### v3

@@ 1 Introduction
Rabiner introduces HMMs [16] and proposes to use the Forward-Backward algorithm [2] for inference. Zhang et al. distill a language model into an HMM [20]. HMMs can be extended to Dynamic Bayesian Networks (DBNs) [7], which can model interactions between different state variables. Inference is done by, e.g, the interface algorithm [14]. However, query answering in Bayesian networks [15] is NP-hard [6], giving rise to approximations. Boyen and Koller represent the distribution over the random variables with successors in the next time step as a product of marginals [5], which is extended by Murphy and Weiss to always use a factored representation [13]. Instead of instantiating a DBN for each time step, Gao et al. introduce a sliding window approach [8]. Murphy provides an overview over approximate inference [14]. Though, the focus in literature is on approximate inference in DBNs or approximation for learning HMMs, but not on approximate inference within HMMs, which we propose in this paper. Vithanage et al. [19] effectively use a top-$k$ approach, i.e., the $k$ most probable transitions, and show that this minimizes the Kullback-Leibler divergence if restricting to $k$ transitions. Contrary, we provide a more fine-grained error assertion. Also, when fixing $p$, one indirectly fixes $k$, too.
@@ 2 Hidden Markov Models Generalizing Dynamic Models
A Hidden Markov Model consists of two series of random variables $(S_{t},O_{t})_{t}$ over a set $S$ of states and $O$ of observations, where $S_{t}$ denotes the state random variable in time step $t$ and $O_{t}$ the respective observation random variable. The probability $P(O_{t}\mid S_{t})$ of an observation depends only on the current state. The probability distribution over the states is given as the temporal behavior $P(S_{t}\mid S_{t-1})$ and a prior $P(S_{0})$. The semantics for a time step $t$ is thus given by
@@
where $P(S_{t})$ is the distribution specifying the probabilities $s_{t}$ for each state $s\in S$.
@@
Now, the important part is that the sum in Equation 1 goes over all possible states, rendering the equation inefficient for large state spaces, like in language models. Our goal is to reduce the number of states that have to be enumerated. While an HMM seems simple, it can be thought of as a generalized dynamic PGM: We have a, potentially complex, state $S_{t}$ emitting observations $O_{t}$. Let us denote the distributions $P(s_{t}),\,s_{t}\in S,$ as forward message updating the state probabilities through time. DBNs [7] can be viewed as HMMs by treating the joint probability distribution over the interface [14] as state and all other random variables jointly as observation. Furthermore, GPT-based language models can be viewed as HMMs and learn the forward message in a highly complex state space. At their core, they can be viewed as HMMs, interpreting, analogous to DBNs, each possible complex internal state of the language model as a state of an HMM with respective transition and observation (or emission) probabilities. Generating sentences from language models is then the same as predicting states and sampling observations from HMMs, which leads to a noisy distribution in the far future without any observations.
@@ 3 Inference in Top-$p$ Hidden Markov Models
In other words, events with high probabilities are promoted while events with low probabilities are ignored, effectively amplifying the differences. For Example 2 and $p=0.9$, Table 2 shows the top-$p$ distribution. As an HMM consists of conditional probability distributions, we independently apply the top-$p$ distributions to them:
@@
Given an HMM $H$ with a set $S$ of states, a set $O$ of observations, probability distributions over the states given by $P_{H}(S_{0})$ and $P_{H}(S_{i+1}\mid s_{i})$, for each state $s_{i}\in S$, and probability distributions $P_{H}(O_{i}\mid s_{i})$, $s_{i}\in S$, over the observations, the top-$p$ HMM of $H$ is another HMM $Q$ over the same set $S$ of states and same set $O$ of observations. The probability distributions of $Q$ are the top-$p$ distributions of the respective ones of $H$, that is, (i) $P_{Q}(S_{0})$is the top-$p$ distribution of $P_{H}(S_{0})$, (ii) $P_{Q}(S_{i+1}\mid s_{i})$is the top-$p$ distribution of $P_{Q}(S_{i+1}\mid s_{i})$, for each state $s_{i}\in S$, and (iii) $P_{Q}(O_{i}\mid s_{i})$is the top-$p$ distribution of $P_{H}(O_{i}\mid s_{i})$, $s_{i}\in S$.
@@
Having shifted the probability mass, we now exploit the new distributions for efficient inference without enumerating all states. Since the transition matrix still has the same dimensions, but is more spare, we use sparse matrices to only enumerate the non-zero entries. We explain how to perform inference, by (sparse) matrix multiplication, in a top-$p$ HMM $Q$ of an original HMM: (1) The top-$p$ HMM $Q$ is built according to Definition 4 by using Algorithm 1. (2) The obtained transition probabilities and observation probabilities are stored in a transition and observation matrix, respectively. (3) The two matrices are each transformed into compressed row storage [1]: For each row in the matrix, a tuple $(c,x)$ for each non-zero entry, where $c$ is the column number and $x$ the value. (4) For inference, sparse matrix-vector multiplication [4] is used to proceed the current state distribution in time and to get the observation distribution: The dot product between each row and the vector is calculated by summing the products of $x$ and the entry $c$ of the vector for each $(c,x)$ tuple in the row. Sparse matrix-vector multiplication runs in time proportional to the number of non-zero entries [4] and is also available for GPUs [3].
@@ 4.2 Overall Approximation Error
Fixing two state distributions $\varphi$ and $\psi$, the transition of a stochastic process $Q$ can be split into two steps: First, a transition defined by $R^{\Gamma}$ from $\Omega$ to $\tilde{\Omega}$ and, second, a transition defined by $R^{\Delta}$ from $\tilde{\Omega}$ to $\Omega$, where $\Omega=\left\{\omega_{i}\right\}_{i}$ is the state space of $Q$ and $\tilde{\Omega}=\Omega\cup\{c\}$ with a new intermediate state $c$. The process $R^{\Gamma}$ preserves its state with probability $1-\gamma$ and transitions to state $c$ with probability $\gamma$. The process $R^{\Delta}$ transitions from $\omega_{i}$ to $\omega_{j}$ with probability $\frac{Q_{i,j}^{\Delta}}{1-\gamma}$. For state $c$, $R^{\Delta}$ transitions to state $\omega_{j}$ with probability $\sum_{i}\varphi(\omega_{i})\frac{Q_{i,j}^{\Gamma}}{\gamma}$. The matrices $Q^{\Gamma}$ and $Q^{\Delta}$ arise from an additive contraction decomposition $Q=Q^{\Gamma}+Q^{\Delta}$, where $Q^{\Gamma}$ and $Q^{\Delta}$ are non-negative matrices such that, for all $i$, $\sum_{j}Q_{i,j}^{\Gamma}=\gamma$, and for all $j$, $\sum_{i}\varphi(\omega_{i})Q_{i,j}^{\Gamma}=\sum_{i}\psi(\omega_{i})Q_{i,j}^{\Gamma}$.
@@
$\delta(\varphi^{\prime},\psi^{\prime})\leq\left(1-\gamma\right)\delta(\varphi,\psi).$ (14)
@@ 5 Evaluation
We evaluate the top-$p$ approach on three HMMs: One favoring our approach, one that is disadvantageous for the top-$p$ approach, and a language model for a realistic application. We investigate the sparsity, the runtime with and without observations, and the total variation distance of the top-$p$ model. We evaluate top-$p$ for the $p$ values $0.5$, $0.7$, and $0.9$ and for the maximum time step $50$. In the runtime evaluation, we solely report numbers for $p=0,9$, as the three top-$p$ approaches are mostly comparable. For the error investigation, we also report numbers for the other tested $p$ values. We run all tests on a 13th Gen Intel(R) Core(TM) i5-1345U CPU with 1.60 GHz and 16 GB of RAM. We use NumPy [10] and SciPy [18] for matrix and sparse matrix operations, respectively.
For the different HMMs, we have trained a very simple language model (LM HMM) as an HMM with 7620 states. In practice, one would distill a larger model into an HMM [20]. However, in this evaluation, we focus on the runtime and not on the quality of the language model. We create a synthetic Bell HMM, which is designed to favor the top-$p$ approach. The Bell HMM distributes, in each next-state distribution, a probability mass of 0.9 over five states and the remaining 0.1 over 795 states, favoring our approach by having lots of unlikely successor states. For a synthetic HMM disadvantageous for the top-$p$ approach, we create a Uniform HMM consisting of uniform state distributions for a system of 800 states. The number of observations equals the number of states in all HMMs.
@@
Sparsity: We define the sparsity as the ratio of zero-entries in the transition matrix. In the Bell HMM, all three top-$p$ approaches lead to a sparsity of above $0.99$. In the Uniform HMM, the sparsity is approximately 1p1-p. And in the LM HMM, the sparsity is for all three top-$p$ greater than $0.97$.
@@
Runtime for Inference: The top-$0.9$ HMM for the Bell HMM outperforms exact inference in the original HMM by a factor of more than $15$. Moreover, all three top-$p$ approaches are significantly faster than inference in the original HMM and are comparable fast. For the Uniform HMM, top-$0.9$ is around $3.39$ times slower than the original HMM. Clearly, the lower $p$, the faster the inference. The reason for the slower runtime is the low sparsity: Sparse matrix multiplication only pays off for high sparsity, which increases with higher $p$. In the LM HMM, top-$p$ is more than $18$ times faster than the original HMM. Figure 1 shows the experiments for the three tested HMMs.
(a) Bell HMM
(b) Uniform HMM
(c) LM HMM
@@
Figure 1: Runtime for inference in the test HMMs. The blue line shows the runtime in the original HMM and the other three ones for different top-$p$ ones.
Runtime for Filtering: Before, we reported runtimes without observations. Now, we enter an observation every five time steps. We omit the LM HMM, since observation in text generation is not useful. In the Bell HMM, the top-$0.9$ approach is more than $77$ times faster than filtering in the original HMM. In the Uniform HMM, in contrast to regular prediction, the top-$0.9$ approach is now more than $24$ times faster than the original HMM, due to more probabilistic calculations helping the overhead induced by sparse matrices to pay off. Figure 2 shows the experiment for the two tested HMMs.
(a) Bell HMM with filtering
(b) Uniform HMM with filtering
Figure 2: Runtime for filtering in two test HMMs. The blue line shows the runtime in the original HMM and the other three ones for different top-$p$ ones.
Total Variation Distance: The total variation in the Bell HMM is $0.086$ for top-$0.9$, $0.285$ for top-$0.7$ and $0.78$ for top-$0.5$. In the Uniform HMM, the total variation distance is $0.099$, $0.299$ and $0.499$ for top-$0.9$, top-$0.7$ and top-$0.5$, respectively. In the LM HMM, the total variation distance is $0.059$, $0.172$ and $0.3$ for top-$0.9$, top-$0.7$ and top-$0.5$, respectively.
Total Variation Distance for Filtering: With observations, the total variation in the Bell HMM is $0.095$ for top-$0.9$, $0.491$ for top-$0.7$ and $0.846$ for top-$0.5$. The increased error compared to without filtering is because more approximations are made in the process. In the Uniform HMM, the total variation distance is the same as for no observations, since the model is uniform. We again omit the LM HMM, since observation in text generation is not useful.
@@
Regarding which $p$ to choose, our evaluation suggests to start with higher values like $0.9$ and adapt them based on the requirements for runtime and total variation. Summing up, the top-$p$ approach significantly increases sparsity to more than $0.9$ for top-$0.9$. Moreover, top-$p$ is at least an order of magnitude faster than the original HMMs, while the total variation distance for top-$0.9$ stays below $0.09$.